OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3) A. y equals negative two-thirds x plus 3 B. y equals negative two-thirds x plus five-thirds C. y = 2x – 21 D. y equals negative two-thirds x minus 9

OpenStudy (anonymous):

So first you want to solve the equation for y.

OpenStudy (anonymous):

3y = 5 - 2x. y = -2/3x + 5/3

OpenStudy (anonymous):

m = -2/3x.

OpenStudy (anonymous):

And then plug in the point-slope form.

OpenStudy (anonymous):

so B

OpenStudy (anonymous):

No A

OpenStudy (anonymous):

hmm I must of did something wrong

OpenStudy (anonymous):

Equation should be y + 3 = -2/3(x - 9)

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

Did you get it now